OpenStudy (moongazer):

help me convert to radian -5rad

OpenStudy (moongazer):

is it 900?

OpenStudy (anonymous):

do you mean cinvert to degrees?

OpenStudy (moongazer):

it is said convert to radians

OpenStudy (moongazer):

the answer is – 900°/π ≈ - 286.48° and i don't know why

OpenStudy (anonymous):

to convert radians to degrees you multiply by 180/pi

OpenStudy (anonymous):

pi radians = 180 degrees

OpenStudy (anonymous):

-5 * 180/pi = -900/pi degrees

OpenStudy (moongazer):

thanks

OpenStudy (moongazer):

so it is radian?

OpenStudy (anonymous):

yw

OpenStudy (moongazer):

-900/pi degrees

OpenStudy (moongazer):

is that a radian?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

=900 / pi = -286.5 degrees

OpenStudy (anonymous):

-5 is the radian

OpenStudy (anonymous):

pi radians is the same as 180 degrees - 'half a circle'

OpenStudy (anonymous):

the question must have been convert -5 radians to degrees

OpenStudy (anonymous):

ok?

OpenStudy (moongazer):

ok

OpenStudy (moongazer):

thanks ^_^